Sulfuric acid is needed in in a lead storage battery. A 25.00 mL sample of “battery acid” taken from the functioning battery requires 53.7 mL of 4.552 M NaOH for complete neutralization. What is the molarity of the battery acid?


1
Expert's answer
2020-10-09T13:46:53-0400

Molarity of the battery acid = CM(H2SO4) = unknown

Volume of battery acid = V(H2SO4) = 25.00 mL

Molarity of sodium hydroxide = CM(NaOH) = 4.552 M

Volume of sodium hydroxide = V(NaOH) = 53.70 mL


Solution:

The reaction between sulfuric acid (H2SO4) and sodium hydroxide (NaOH) is:

H2SO4 + 2NaOH = Na2SO4 + 2H2O

According to the equation: n(H2SO4) = n(NaOH)/2

Therefore,

2 × CM(H2SO4) × V(H2SO4) = CM(NaOH) × V(NaOH)

2 × CM(H2SO4) × (25.00 mL) = (4.552 M) × (53.70 mL)

CM(H2SO4) = (4.552 M × 53.70 mL) / (2 × 25.00 mL) = 4.8888 M = 4.889 M

Molarity of the battery acid = CM(H2SO4) = 4.889 M.


Answer: The molarity of the battery acid is 4.889 M.